OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Numerical Answers Expected!

OpenStudy (jazzyfa30):

so lets set it up like they did c=???

OpenStudy (anonymous):

This question is not formed correctly.

OpenStudy (jazzyfa30):

@Mikeyy1992 ur not suppose to give them the answer your suppose to help them get the answer

OpenStudy (anonymous):

As you can see I deleted it, because the question was formed incorrectly.

OpenStudy (anonymous):

but the thing was i got i wrong before and i used 76.50

OpenStudy (ivettef365):

substitute the 100 for x so you have c= 0.765 (100) + 0.06 [(0.765 (100)]

OpenStudy (ivettef365):

so you have C = 76.50 + 4.59 C= 81.09

OpenStudy (ivettef365):

81.09 is the price with tax

OpenStudy (anonymous):

I think after you put in the substitute and solve and get 76.50 you put 15% of that and subtract then subtract 10% of that. I got 81.09 on the first time he asked this question, but the 15% and 10% confused me.

OpenStudy (anonymous):

When he didn't include the percents, i got 81.09.

OpenStudy (ivettef365):

if you get 100 and with the discount of 15% = 85.00 then $85 with the discount of 10% = 76.50 but when you add tax of .06 then total to pay = 81.09

OpenStudy (anonymous):

Well perfect I was correct Lol.

OpenStudy (anonymous):

No medal for me *cries*